OpenStudy (anonymous):

Solve by substitution y = 2x + 7 4x − 3y = −9

OpenStudy (anonymous):

4x - 3(2x+7) = 9 since if y = 2x+7 you can simply replace the y in the 2nd equation with what y is in the first then rules of solving for x 4x-6x-21=9 -2x=30 x=-15